please help me solve this question: Determine if line y=6x=12 intersects the ellipse x^2/25=y^2/9=1.give the coordinates of the point of intersection if applicable
Answer by Alan3354(69443)

please help me solve this question: Determine if line y=6x=12 intersects the ellipse x^2/25=y^2/9=1.give the coordinates of the point of intersection if applicable
----------------
y=6x=12 Too many equal signs